### v1

# Comment on “Galvano-rotational effect induced by electroweak interactions in pulsars”

Abstract
In this comment, we obtain the corrected energy spectrum for Dvornikov’s paper, that is, a spectrum in which the matter term, given by $V_{L}$, is “tied” in both the radial and angular quantum numbers, given by $N$ and $J_{z}$. In particular, the error committed by Dvornikov arose from an incorrectly solved second-order differential equation. Besides, such an equation also had another error, i.e. since Dvornikov considered ultrarelativistic particles where $m\to 0$, it implies that all terms of this equation containing $m$ should vanish; however, that is not what happened. Therefore, solving in detail the corrected Dvornikov equation, we obtain the corrected spectrum, where we verified that for $J_{z}>0$ (positive angular momentum), $\sigma=+1$ (upper component of the spinor) and $V_{L}>0$, or $J_{z}<0$ (negative angular momentum), $\sigma=-1$ (lower component of the spinor) and $V_{L}<0$, we recover the particular spectrum of Dvornikov’s paper with $m^{2}=0$ (as it should be).
I Introduction
In a paper published in the Journal of Cosmology and Astroparticle Physics (JCAP), entitled “Galvano-rotational effect induced by electroweak interactions in pulsars”, Dvornikov [2] studied the electroweakly interacting particles in rotating matter, where the existence of the electric current along the axis of the matter rotation was predicted in this system. According to him, this new Galvano-rotational effect is caused by the parity-violating interaction between massless charged particles in the rotating matter. To do such a study, he started with the exact solution of the Dirac equation for a fermion involved in the electroweak interaction in the rotating frame with constant angular velocity in cylindrical coordinates, where he used the vierbein vectors formalism to write the Dirac equation in a curved spacetime. Then, using the obtained solution (for $m\to 0$, i.e. ultrarelativistic neutrinos), he derived the induced electric current which turns out to flow along the rotation axis, and studied the possibility of the appearance of the Galvano-rotational effect in the dense matter of compact astrophysical objects. Later, he discussed the particular example of neutron and hypothetical quark stars, where it was shown that using this effect, one can expect the generation of toroidal magnetic fields comparable with poloidal ones in old millisecond pulsars. Besides, he also briefly discussed the generation of the magnetic helicity in these stars. Finally, he analyzed the possibility of applying the Galvano-rotational effect for the description of the asymmetric neutrino emission from a neutron star to explain pulsar kicks. In particular, this paper is well-written and addresses a very interesting topic involving Dirac neutrinos in rotating frames of reference and interacting with a background matter. Besides, the formalism used by him is also very important in the literature when working with Dirac fermions in curved space-times or rotating frames [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
However, analyzing in detail the energy spectrum in Dvornikov’s paper [2], the following question/doubt arose: Why does only a part of the Dvornikov energy spectrum depend on the angular quantum number $J_{z}$? (in which is given by $J_{z}\omega$, where $\omega$ is the angular velocity of the rotating frame). In other words, Why is the matter term in which the neutrino interacts, given by $V_{L}$, not also “tied” to the angular quantum number $J_{z}$? (in which is given by $4|V_{L}|N\omega$, where $N$ is the radial quantum number). To be more specific, the matter term in Dvornikov’s paper [2] that we refer to is that found inside a square root (which is a feature of any relativistic spectrum) and not that found outside, in which it is a free term (i.e. a term “tied” to nothing). In particular, such a question/doubt arose from the fact that Dvornikov [2] did not make it explicit/clear why $J_{z}$ also did not appear inside of the square root (or “tied” to $V_{L}$), such as occurs with $N$. In this way, this motivated us to look by the correct spectrum for the neutrino, that is, a spectrum in which the matter term is “tied” in both $N$ and $J_{z}$, and not only to $N$ as shown by Dvornikov [2]. Therefore, the goal of the present comment is to obtain the corrected energy spectrum for Dvornikov’s paper [2]. To achieve this, we use as a basis Refs. [3, 4, 7, 9, 10, 11, 12, 13, 14, 16, 17], where the Dirac equation was worked in curved space-times (with and without rotation). Besides, it is also important to mention that something similar to this has also been done by Dvornikov in another paper of his, but published in the JHEP [18] (i.e. the energy spectrum did not depend on $J_{z}$, whether inside or outside the square root). However, recently the corrected/complete energy spectrum of this paper was obtained by Ref. [19].
II Quick review of the main steps that Dvornikov took to obtain the energy spectrum
According to Dvornikov [2], the Dirac equation for a massive neutrino (of mass $m$) involved in the parity-violating interaction and moving in a (3+1)D curved space-time has the form [2]:
$[i\gamma^{\mu}(x)\nabla_{\mu}-m]\psi=\gamma_{\mu}(x)\left\{\frac{V^{\mu}_{L}}{2}[1-\gamma^{5}(x)]+\frac{V^{\mu}_{R}}{2}[1+\gamma^{5}(x)]\right\}\psi,$ (1)
where $\gamma^{\mu}(x)=e^{\ \mu}_{a}\gamma^{a}$ and $\gamma_{\mu}(x)=e^{a}_{\ \mu}\gamma_{a}$ are the curved gamma matrices, $\gamma^{a}$ and $\gamma_{a}$ are the flat gamma matrices, $e^{\ \mu}_{a}$ is the vierbein and $e^{a}_{\ \mu}$ is its inverse (and satisfy the orthogonality condition $e^{a}_{\ \mu}e^{\ \mu}_{b}=\delta^{a}_{b}$), $\nabla_{\mu}=\partial_{\mu}+\Gamma_{\mu}$ is the covariant derivative, $\Gamma_{\mu}=-\frac{i}{4}\sigma^{ab}\omega_{ab\mu}=-\frac{i}{4}\sigma^{ab}e^{\ \nu}_{a}e_{b\nu;\mu}$ is the spin connection, $\sigma^{ab}=\frac{i}{2}[\gamma^{a},\gamma^{b}]$ are the generators of the Lorentz transformations in a locally Minkowskian frame, $\gamma^{5}(x)=-\frac{i}{4!}E^{\mu\nu\alpha\beta}\gamma_{\mu}(x)\gamma_{\nu}(x)\gamma_{\alpha}(x)\gamma_{\beta}(x)$ is the curved fifth gamma matrix, $E^{\mu\nu\alpha\beta}=\frac{1}{\sqrt{-g}}\varepsilon^{\mu\nu\alpha\beta}$ is the covariant antisymmetric tensor in curved space-time, $g$ is the determinant of the metric (i.e. $g=$ det$(g_{\mu\nu})$), $V^{\mu}_{L,R}$ are the effective potentials (with $V^{\mu}_{L}\neq V^{\mu}_{R}$), i.e. the last term of Eq. (1) is the background matter or simply the matter term, and $\psi$ is the four-component Dirac spinor, respectively.
According to Dvornikov [2], the line element in the rotating frame is written as:
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=(1-\omega^{2}r^{2})dt^{2}-dr^{2}-2\omega r^{2}dtd\phi-r^{2}d\phi^{2}-dz^{2},$ (2)
where $x^{\mu}=(t,r,\phi,z)$ is the position four-vector in cylindrical coordinates, $\omega$ is the constant angular velocity, and $g_{\mu\nu}$ is the metric tensor, respectively. In addition, one can check that the metric tensor in Eq. (2) can be diagonalized (via $\eta_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}g_{\mu\nu}$) if we use the following vierbein vectors [2]:
$e_{0}^{\ \mu}=\left(1,0,-\omega,0\right),$
$e_{1}^{\ \mu}=\left(0,1,0,0\right),$
$e_{2}^{\ \mu}=\left(0,0,\frac{1}{r},0\right),$
$e_{3}^{\ \mu}=(0,0,0,1),$ (3)
where $\eta_{ab}$=diag(+1,1,1,1)$1$(+1,-1,-1,-1)( + 1 , - 1 , - 1 , - 1 ) is the metric in a locally Minkowskian frame. With this, the only non-zero component of the connection one-form $\omega_{ab}=\omega_{ab\mu}dx^{\mu}$ is given as follows [2] (originally obtained by Ref. [3]):
$\omega_{12\mu}=-\omega_{21\mu}=(\omega,0,1,0),$ (4)
where implies that $i\gamma^{\mu}(x)\Gamma_{\mu}=\frac{i}{2r}\gamma^{1}$ [2, 3].
From this, Dvornikov [2] rewrote Eq. (1) in the following equation (with $V^{\mu}_{L,R}=(V^{0}_{L,R},0,0,0)=(V_{L,R},0,0,0)$ and $\gamma^{5}(x)=\gamma^{5}=\gamma_{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$):
$\mathcal{D}\psi=(\gamma^{0}-\omega r\gamma^{2})(V_{V}-V_{A}\gamma^{5})\psi,$ (5)
where
$\mathcal{D}=\left[i\gamma^{0}(\partial_{0}-\omega\partial_{\phi})+i\gamma^{1}\left(\partial_{r}+\frac{1}{2r}\right)+i\gamma^{2}\frac{\partial_{\phi}}{r}+i\gamma^{3}\partial_{z}-m\right],$ (6)
being $V_{V,A}=(V_{L}\pm V_{R})/2$ the vector and axial parts of the effective potentials, and the matrices $\gamma^{0}$, $\gamma^{k}$ $(k=1,2,3)$, and $\gamma^{5}$ are written as (chiral representation):
[…]
So, considering the following spinor:
$\psi=\text{exp}(-iEt+iJ_{z}\phi+ip_{z}z)\psi_{r},$ (8)
Dvornikov [2] obtained:
$\left[\gamma^{a}Q_{a}-m+V\right]\psi_{r}=0,$ (9)
where $J_{z}=\pm\frac{1}{2},\pm\frac{3}{2},\pm\frac{5}{2},\ldots$ is the angular quantum number (i.e. a quantum number associated with the angular part of the spinor), $Q^{a}=q^{a}-q_{\text{eff}}A^{a}_{\text{eff}}$, being $q_{\text{eff}}$ the effective electric charge, $q^{a}=\left(E+J_{z}\omega-V_{V},-i\partial_{r},0,p_{z}\right)$, $V=V_{A}(\gamma^{0}-\omega r\gamma^{2})\gamma^{5}$, and $A^{a}_{\text{eff}}$ is the potential of the effective electromagnetic field, given as follows:
$A^{a}_{\text{eff}}=\left(0,\frac{i}{2q_{\text{eff}}r},\frac{1}{q_{\text{eff}}}\left[V_{V}\omega r-\frac{J_{z}}{r}\right],0\right).$ (10)
Besides, Dvornikov [2] considered the solution of Eq. (9) in the form: $\psi_{r}=[\gamma^{a}Q_{a}+m-V]\Phi$, where $\Phi=\Phi(r)$ is a new spinor. With respect to the form of $\psi_{r}$, it aims to arrive at a “quadratic Dirac equation”, that is, arriving at a second-order differential equation without passing directly by the first-order differential equations coupled with the spinor components, such as is done in Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 17]. In particular, this method for obtaining a “quadratic Dirac equation” also has great relevance in the literature, where several papers have already used it [16, 20, 21, 22, 23, 24, 25, 26, 27].
Therefore, using (10) and a the form of $\psi_{r}$ (defined just above), Dvornikov [2] obtained the following second-order differential equation (or “quadratic Dirac equation”) for the spinor $\Phi$:
$\left[\left(\partial_{r}+\frac{1}{2r}\right)^{2}+(E+J_{z}\omega-V_{V})^{2}-\left(\frac{J_{z}}{r}-V_{V}\omega r\right)^{2}+2V_{A}\gamma^{5}\left[\left(\frac{J_{z}}{r}-V_{V}\omega r\right)\omega r-(E+J_{z}\omega-V_{V})+\frac{\omega}{2}\Sigma_{3}\right]\right]\Phi,$
$+\left[\left(V_{V}\omega+\frac{J_{z}}{r^{2}}\right)\Sigma_{3}-p^{2}_{z}+V_{A}^{2}(1-\omega^{2}r^{2})\gamma^{5}+2mV_{A}(\gamma^{0}-\omega r\gamma^{2})\gamma^{5}-m^{2}\right]\Phi=0,$ (11)
where $\Sigma_{3}=\gamma^{0}\gamma^{3}\gamma^{5}$.
So, according to Dvornikov [2], the solution of Eq. (II) can be found for ultrarelativistic particles (i.e. in the limit $m\to 0$). With this, it is possible write $\Phi$ in the form $\Phi=\upsilon\varphi$, where $\upsilon$ is a constant spinor (and satisfies $\Sigma_{3}\upsilon=\sigma\upsilon$ and $\gamma^{5}\upsilon=\chi\upsilon$, where $\sigma=\chi=\pm 1$) and $\varphi=\varphi(r)$ is a scalar function [2]. Besides, Dvornikov [2] first studied left particles where $(1+\gamma^{5})\psi=0$, in which corresponds $\chi=+1$ (the case of right particles with $\chi=-1$ can be studied analogously). In this way, using a new variable in Eq. (II) given by $\rho=|V_{L}|\omega r^{2}$ (i.e. making a change of variable), Dvornikov [2] obtained the following equation for $\varphi_{\sigma}$:
$\left[\rho\partial^{2}_{\rho}+\partial_{\rho}-\frac{1}{4\rho}\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right)^{2}-\frac{\rho}{4}-\frac{1}{2}\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right)+\kappa\right]\varphi_{\sigma}=0,$ (12)
where
$\kappa=\frac{1}{4|V_{L}|\omega}\left[E^{2}+2E(J_{z}\omega-V_{L})-p^{2}_{z}-m^{2}+(V_{L})^{2}+J^{2}_{z}\omega^{2}+V_{L}\omega\sigma\right]+\frac{1}{2}\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right),$ (13)
being $J_{z}=(1/2-l)\text{sgn}(V_{L})$, where $l=0,\pm 1,\pm 2,\ldots$.
According to Dvornikov [2], the solution of Eq. (12) (which vanishes at infinity, i.e. $\varphi_{\sigma}\to 0$ when $r\to\infty$) depends on the associated Laguerre polynomials; consequently, the parameter $\kappa$ must be equal to a positive integer, that is, $\kappa=N$, where $N=0,1,2,\ldots$ is the radial quantum number (i.e. a quantum number associated with the radial/spatial part of the spinor). Therefore, using the condition $\kappa=N$, Dvornikov [2] obtained the following relativistic energy spectrum for the Dirac neutrino (written in terms of $l$):
$E=V_{L}+\left(l-\frac{1}{2}\right)\omega\text{sgn}(V_{L})\pm\mathcal{E},\ \ \mathcal{E}=\sqrt{p^{2}_{z}+m^{2}+4|V_{L}|N\omega}.$ (14)
According to the spectrum above, we clearly see that the matter term (given by $V_{L}$) inside the square root is not “tied” to the angular quantum number $J_{z}$ (or $l$), but only to the radial quantum number $N$ (given by $4|V_{L}|N\omega$). That is, only a part of the Dvornikov energy spectrum depends on $J_{z}$, which is the second term of the spectrum and is “tied” to $\omega$ (given by $J_{z}\omega$). Besides, we also think very strange the presence of the (quadratic) mass term in the spectrum (given by $m^{2}$) since Dvornikov [2] said that the solution of Eq. (II) can be found for ultrarelativistic particles (i.e. $m\to 0$). In other words, this spectrum is incorrect (due to the presence of $m^{2}$ and the absence of $J_{z}$ “tied” to the matter term $V_{L}$ inside the square root).
III The corrected energy spectrum for Dvornikov’s paper
Before we obtain the corrected spectrum for Dvornikov’s paper, we first need to correct Eq. (12). So, as stated by Dvornikov [2], the solution of Eq. (II) can be found for ultrarelativistic particles ($m\to 0$), which implies that all terms of Eq. (12) containing $m$ should vanish (or be zero); however, that is not what happened since there is the presence of $m^{2}$ in the equation. In other words, Dvornikov [2] zeroed one term, given by $2mV_{A}(\gamma^{0}-\omega r\gamma^{2})\gamma^{5}$, and kept another, given by $m^{2}$ (which could not have happened, i.e. it is wrong). Besides, is also important to mention here that the condition for we have only left particles in the Dirac equation (1) is if $V^{\mu}_{R}=0$ (or $V_{R}=0$) [18], and do not for $(1+\gamma^{5})\psi=0$, as stated by Dvornikov [2]. Consequently, this implies that: $V_{V,A}=V_{L}/2\neq 0$ (Dvornikov used this to arrive at (12)). In short, Eq. (12) is only really correct if the quadratic mass term is also zero ($m^{2}=0$); consequently, the equation must not have any terms containing $m$ (in particular, that should be the case since Dvornikov did an earlier paper, given by [18], in which the equation does not contain $m$).
So, making $m^{2}=0$ as well as excluding the terms ±$\pm\frac{1}{2}\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right)$ in Eq. (12), that is, there is no need to use such terms, we have the following equation (with $\partial_{\rho}\to\frac{d}{d\rho}$):
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\frac{d}{d\rho}-\frac{\gamma^{2}}{4\rho}-\frac{\rho}{4}+\beta\right]\varphi_{\sigma}(\rho)=0,$ (15)
where we define:
$\gamma=\gamma_{\sigma}\equiv\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right),\ \ \beta=\beta_{\sigma}\equiv\frac{1}{4|V_{L}|\omega}\left[E^{2}+2E(J_{z}\omega-V_{L})+(V_{L})^{2}+J^{2}_{z}\omega^{2}+V_{L}\omega\sigma-p^{2}_{z}\right],$ (16)
or better:
$\beta\equiv\frac{1}{4|V_{L}|\omega}\left[(E-V_{L}+J_{z}\omega)^{2}+2\omega V_{L}\left(J_{z}+\frac{\sigma}{2}\right)-p^{2}_{z}\right].$ (17)
Now, let us solve Eq. (15) through of the asymptotic behavior/limit for $\rho\to 0$ and $\rho\to\infty$ (i.e. two asymptotic behaviors) [3, 4, 7, 9, 10, 11, 12, 13, 14, 16, 17]. Therefore, for $\rho\to 0$, we have:
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\frac{d}{d\rho}-\frac{\gamma^{2}}{4\rho}\right]\varphi_{\sigma}(\rho)\approx 0,$ (18)
whose solution is given by $\varphi_{\sigma}(\rho)=\rho^{|\gamma|/2}$.
Now, for $\rho\to\infty$, we have:
$\left[\rho\frac{d^{2}}{d\rho^{2}}-\frac{\rho}{4}\right]\varphi_{\sigma}(\rho)\approx 0,$ (19)
whose solution is given by $\varphi_{\sigma}(\rho)=e^{-\rho/2}$.
In this way, a regular solution for Eq. (15) can be written as:
$\varphi_{\sigma}(\rho)=Ce^{-\rho/2}\rho^{|\gamma|/2}F_{\sigma}(\rho),$ (20)
where $C$ is a normalization constant and $F_{\sigma}(\rho)$ is a unknown function to be determined. In particular, to be a physically allowed solution (or a normalizable solution), $\varphi_{\sigma}(\rho)$ must satisfy two boundary conditions of the system, i.e. vanish at the origin and infinity, such as $\varphi_{\sigma}(\rho\to 0)=\varphi_{\sigma}(\rho\to\infty)=0$.
So, replacing the solution (20) in (15), we obtain a second-order differential equation for $F_{\sigma}(\rho)$:
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\left(\left|\gamma\right|+1-\rho\right)\frac{d}{d\rho}-\left(\frac{\left|\gamma\right|+1}{2}-\beta\right)\right]F_{\sigma}(\rho)=0,$ (21)
or yet (“Dvornikov-type form”)
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\left(|\gamma|+1-\rho\right)\frac{d}{d\rho}+\kappa\right]F_{\sigma}(\rho)=0,\ \ \kappa\equiv\left(-\frac{|\gamma|+1}{2}+\beta\right).$ (22)
According to Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], the equation above is an associated Laguerre equation (or a confluent hypergeometric equation), whose solution are the associated Laguerre polynomials (or confluent hypergeometric functions). Consequently, the parameter $\kappa$ must be equal to a positive integer, that is, $\kappa=N$, where $N=0,1,2,\ldots$ is the radial quantum number (i.e. a quantum number associated with the radial/spatial part of the spinor). In fact, the correct thing to say would be that the last term of (21) is a negative integer, but since the minus sign has already been incorporated into $\kappa$, it amounts to the same thing/result. Therefore, from the condition $\kappa=N$ (a quantization condition), we obtain the following corrected energy spectrum for Dvornikov’s paper:
$E=V_{L}-J_{z}\omega\pm\mathcal{E},\ \ \mathcal{E}=\sqrt{p^{2}_{z}+4|V_{L}|N_{\sigma}\omega},\ \ (J_{z}=(1/2-l)\text{sgn}(V_{L})),$ (23)
where we define
$N_{\sigma}\equiv\left(N+\frac{1}{2}+\frac{|J_{z}-\frac{\sigma}{2}|}{2}-\textbf{sgn}(V_{L})\frac{(J_{z}+\frac{\sigma}{2})}{2}\right)=\left(N+\frac{1-\sigma\text{sgn}(V_{L})}{2}+\frac{|J_{z}-\frac{\sigma}{2}|}{2}-\text{sgn}(V_{L})\frac{(J_{z}-\frac{\sigma}{2})}{2}\right),$ (24)
being $N_{\sigma}=N_{eff}=N_{total}$ a type of “effective or total quantum number” (since it depends on all others), and is similar to that of Refs. [10, 11, 12, 13, 14, 28, 29], and we use $V_{L}=\text{sgn}(V_{L})|V_{L}|$. In Table 1, we have the values of $N_{\sigma}$ for eight possible combinations of $\sigma$, $\text{sgn}(V_{L})$, and $J_{z}$. So, we see clearly/explicitly that for $J_{z}>0$ (neutrino with positive angular momentum), $\sigma=+1$ (upper component of the Dirac spinor) and $V_{L}>0$, or $J_{z}<0$ (neutrino with negative angular momentum), $\sigma=-1$ (lower component of the Dirac spinor) and $V_{L}<0$, we recover the particular energy spectrum of Dvornikov’s paper with $m^{2}=0$ (as it should be) [2].
Table 1: Values of $N_{\sigma}$ for eight possible combinations of $\sigma$, $\text{sgn}(V_{L})$, and $J_{z}$.
$N_{\sigma}$ $\sigma$ $\text{sgn}(V_{L})$ $J_{z}$
$N_{\sigma}=N$ $\sigma=+1$ $\text{sgn}(V_{L})=+1$ $J_{z}>0$
$N_{\sigma}=N+\frac{1}{2}-J_{z}$ $\sigma=+1$ $\text{sgn}(V_{L})=+1$ $J_{z}<0$
$N_{\sigma}=N+\frac{1}{2}+J_{z}$ $\sigma=+1$ $\text{sgn}(V_{L})=-1$ $J_{z}>0$
$N_{\sigma}=N+1$ $\sigma=+1$ $\text{sgn}(V_{L})=-1$ $J_{z}<0$
$N_{\sigma}=N+1$ $\sigma=-1$ $\text{sgn}(V_{L})=+1$ $J_{z}>0$
$N_{\sigma}=N+\frac{1}{2}-J_{z}$ $\sigma=-1$ $\text{sgn}(V_{L})=+1$ $J_{z}<0$
$N_{\sigma}=N+\frac{1}{2}+J_{z}$ $\sigma=-1$ $\text{sgn}(V_{L})=-1$ $J_{z}>0$
$N_{\sigma}=N$ $\sigma=-1$ $\text{sgn}(V_{L})=-1$ $J_{z}<0$
IV Final remarks
In the present comment, we obtain the corrected energy spectrum for Dvornikov’s paper [2], that is, a spectrum in which the matter term, given by $V_{L}$, is “tied” in both the radial and angular quantum numbers, given by $N$ and $J_{z}$ (where $N\geq 0$ and $J_{z}\gtrless 0$), and not only to $N$ as shown by Dvornikov. To be more specific, the matter term in Dvornikov’s paper [2] that we refer to is that found inside a square root and not that found outside, in which it is a free term. In particular, the error committed by Dvornikov arose from an incorrectly solved second-order differential equation (after having made a change of variable), where only one asymptotic behavior was considered, in which the scalar function vanishes at infinity (the correct is to vanish at both origin and infinity). Besides, such differential equation also had another error, i.e. since Dvornikov considered ultrarelativistic particles where $m\to 0$ (“massless neutrinos”), it implies that all terms of this equation containing $m$ should vanish (or be zero); however, he zeroed one term and kept another (which could not have happened). Therefore, using as a base several papers in the literature on the Dirac equation (in curved space-times, with and without rotation), we solve in detail the corrected Dvornikov equation and obtain the corrected spectrum for the Dirac neutrinos. Finally, we verified that for $J_{z}>0$ (positive angular momentum) with $\sigma=+1$ (upper component of the Dirac spinor) and $V_{L}>0$, or $J_{z}<0$ (negative angular momentum) with $\sigma=-1$ (lower component of the Dirac spinor) and $V_{L}<0$, we recover the particular energy spectrum of Dvornikov’s paper with $m^{2}=0$ (as it should be).
Acknowledgments
The author would like to thank the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) for financial support through the postdoc grant No. 175392/2023-4, and also to the Department of Physics at the Universidade Federal da Paraíba (UFPB) for hospitality and support.
Data availability statement
This manuscript has no associated data or the data will not be deposited. [Author’ comment: There is no data associated with this manuscript or no data has been used to prepare it.]